# FGBERT: Function-Driven Pre-trained Gene Language Model for Metagenomics

### v1

@@ 4.6 Ablation Study
Figure 5: Ablation studies of proposed modules on four downstream tasks.
[…]
@@
Figure 6: t-SNE visualization of input protein embeddings and our proposed protein-based gene embeddings for ATP synthase sequences (n=1177). Each icon denotes an ATP synthase sequence and is colored according to different functions.
Figure 7: Comparative Analysis on Tokenization Efficiency: Balancing Time(s) vs. Memory Usage(MB). Each circle point on the plot corresponds to a specific dataset, with the point size indicating the dataset’s scale.
[…]
@@
Furthermore, we perform a visualization experiment to validate the effectiveness of MGM in One-to-Many scenario. ATP synthases can exhibit different functions in different organisms to adapt to their respective environmental conditions, even though the basic functions are the same [22]. To explore this, we collect 1177 ATP synthase sequences from UniProt [1] and color them according to six taxonomies, i.e. Hydrolases (35.6%, dark green), Isomerases (3.6%, orange), Ligases (35.7%, blue), Lyases (10.7%, pink), Oxidoreduct (8.9%, light green) and Transferases (5.3%, yellow) (numbers in parentheses indicate category ratios). Figure 6a shows the clustering results of ATPase protein embeddings without genome contextual analysis, indicating that more dispersed clustering results of genes in different genome contexts. Figure 6b presents the clustering results of ATPase embedding after our designed protein-based gene embedding, which shows genes belonging to the same category (in the same color) apparently cluster together. Concretely, Isomerases (3.5%, orange) and Transferases (5.3%, yellow), which account for the smallest percentage, cluster together, whereas, in the left plot, these two are scattered. This demonstrates that our proposed MGM can resolve the One-to-Many relationship between sequences and functions.
4.7 Model Efficiency Study
@@
We analyze the time complexity and memory efficiency of token processing for ours and four genome pre-trained methods on six datasets in Figure 7. Our tokenizer achieves the best balance between time, computation, and performance (M.F1) on all datasets. Notably, for NCyc dataset (brown circle) with 213,501 sequences, ours exhibits a remarkable reduction in time by 31.05% compared to DNABERT2 along with a substantial decrease in memory usage by 94.33%. Additionally, for CARD dataset (orange circle) containing 1,966 sequences, ours showcases a significant reduction in time by 61.70% compared to DNABERT2 accompanied by a substantial decrease in memory usage by 58.53%. Although HyenaDNA has a smaller memory footprint than ours on the Operons, CARD, VFDB, and ENZYME datasets, it performs worse than ours in terms of time cost, and performance.
Figure 8: Sensitivity w.r.t Hyper-parameters $\alpha,b$ of CARD dataset on AMR Gene Family.
@@ 4.8 Sensitivity Analysis
Our sensitivity analysis on the batch size $b$ indicates that the proposed model can be optimized effectively using small batches without larger performance degradation shown in Figure 8. This observation is particularly important for resource-constrained scenarios, highlighting our model maintains good performance even with limited data. We select a moderate batch size 1000 that balances computational efficiency and model performance. Next, we analyze another hyper-parameter of balance ratio $\alpha$ on CARD dataset. FGBERT obtains good performance for different values of $\alpha$, demonstrating its robustness and insensitivity to this hyper-parameter. We select $\alpha$ to be 0.4.
@@ 5 Conclusion
In this paper, we propose a new idea of protein-based gene representation, preserving essential biological characteristics within each gene sequence. With the new context-aware tokenizer, we propose MGM, a gene group-level pre-training task, designed to learn the interactions between genes. Additionally, we develop TEM-CL, a contrastive learning module to generate multiple positive and negative samples to distinguish the gene sequences. MGM and TEM-CL constitute a joint pre-training model, FGBERT for metagenomic data. Our experiments and visualizations demonstrate the superior performance of our model. For the future, it remains to be explored how to incorporate multi-omics data, such as metabolomics, into our metagenomic pre-trained model.
Appendix A Prelinaries
[…]
Appendix B Model Configurations
All our experiments are performed on 4 NVIDIA V100 GPUs and the PyTorch framework. The encoder of FGBERT is initialized with Roberta [11]. The parameters for pre-training: batch size = 1000, 19 encoder layers, 10 attention heads, embedding dimension = 1280, and relative position encoding. During the pre-training stage, the model is trained for 500 epochs. We use the AdamW optimizer [35] with a weight decay of 0.02. The learning rate is initialized to 1e-5 and warmed up to 1e-4 after 5000 epochs. Then, it is reduced to 1e-5 by a cosine decay strategy. The overall model comprises 954.73M parameters with a computational load of 2.55B FLOPs.
Appendix C Dataset.
MGnify dataset includes genome sequence information from different microbial communities and their Biome Names (e.g., ”Engineered”, ”Host-associated”, and ”Control”), as well as their corresponding Sample Classes, Sample Numbers, and Spectrum Categories. Lineage Category refers to the taxonomic information of microorganisms, which describes the taxonomic hierarchy of microorganisms from Domain (e.g., ”root” or ”Engineered”) to Species (e.g., ”Continuous culture” or ”Saliva”), which can be observed in Table 5.
Table 5: Microbial Species and Genomic Data Across Various Environments in the MGnify Dataset. Each column is catalogs, species, and genomes respectively.
@@ Appendix B Model Configurations
E.coil K12 RegulonDB Information such as the name of the operon, its description, and the name of the genes included in the partially E. coil K12 RegulonDB dataset are listed in Table 6.
Table 6: Examples of Operons Dataset. Information such as the name of the operon, its description, and the name of the genes included in the partially E. coil K12 RegulonDB dataset
GeneID Operon name
Discription Operon name
@@
2971 ygiV
DNA-binding transcriptional repressor YgiV
None
None
2972 ygiW
BOF family protein YgiW
None
None
2973 qseB
DNA-binding transcriptional activator QseB
qseBC
None
2974 qseC
sensor histidine kinase QseC
qseBC
None
2975 ygiZ
DUF2645 domain-containing inner membrane protein YgiZ
None
None
2976 mdaB
NADPH:quinone oxidoreductase MdaB
None
None
2977 ygiN
putative quinol monooxygenase YgiN
None
None
2978 parE
DNA topoisomerase IV subunit B
nudF-yqiB-cpdA-yqiA-parE
None
2979 yqiA
esterase YqiA
nudF-yqiB-cpdA-yqiA-parE
None
2980 cpdA
cAMP phosphodiesterase
nudF-yqiB-cpdA-yqiA-parE
None
2981 yqiB
DUF1249 domain-containing protein YqiB
nudF-yqiB-cpdA-yqiA-parE
None
2982 nudF
ADP-sugar pyrophosphatase
nudF-yqiB-cpdA-yqiA-parE
None
2983 tolC
outer membrane channel TolC
tolC-ygiBC
tolC-ygiBC
2984 ygiB
DUF1190 domain-containing protein YgiB
tolC-ygiBC
tolC-ygiBC
2985 ygiC
putative acid–amine ligase YgiC
tolC-ygiBC
tolC-ygiBC
2986 ygiD
4,5-DOPA dioxygenase extradiol
None
None
@@
CARD Table 7 shows a partial list of instances from the CARD dataset, including data IDs, ARO (Antibiotic Resistance Ontology) Accession IDs, AMR (Antimicrobial Resistance) category names, drug category names, and Resistance Mechanism names.
@@
Model ID ARO Accession
AMR Gene Family
Drug Class
Resistance Mechanism
2 3002999
CblA beta-lactamase
cephalosporin
antibiotic inactivation
5 3002867
trimethoprim resistant dihydrofolate reductase dfr
diaminopyrimidine antibiotic
antibiotic target replacement
7 3001989
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
10 3002244
CARB beta-lactamase
penam
antibiotic inactivation
20 3002012
CMY beta-lactamase
cephamycin
antibiotic inactivation
24 3003552
Target protecting FusB-type protein conferring resistance to Fusidic acid
fusidane antibiotic
antibiotic target protection
25 3001980
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
27 3002835
lincosamide nucleotidyltransferase (LNU)
lincosamide antibiotic
antibiotic inactivation
31 3003163
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
35 3002804
fosfomycin thiol transferase
phosphonic acid antibiotic
antibiotic inactivation
38 3002649
APH(3’)
aminoglycoside antibiotic
antibiotic inactivation
39 3002528
AAC(3)
aminoglycoside antibiotic
antibiotic inactivation
40 3002771
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
41 3003198
16S rRNA methyltransferase (G1405)
aminoglycoside antibiotic
antibiotic target alteration
46 3002125
CMY beta-lactamase
cephamycin
antibiotic inactivation
50 3002380
SME beta-lactamase
carbapenem
antibiotic inactivation
51 3002538
AAC(3)
aminoglycoside antibiotic
antibiotic inactivation
58 3002762
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
60 3002795
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
63 3002546
AAC(6’)
aminoglycoside antibiotic
antibiotic inactivation
64 3002083
CMY beta-lactamase
cephamycin
antibiotic inactivation
69 3002620
ANT(3”)
aminoglycoside antibiotic
antibiotic inactivation
71 3002778
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
75 3003026
fusidic acid inactivation enzyme
fusidane antibiotic
antibiotic inactivation
88 3002066
CMY beta-lactamase
cephamycin
antibiotic inactivation
92 3001904
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
94 3002092
CMY beta-lactamase
cephamycin
antibiotic inactivation
95 3002067
CMY beta-lactamase
cephamycin
antibiotic inactivation
98 3002059
CMY beta-lactamase
cephamycin
antibiotic inactivation
99 3002753
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
105 3002243
CARB beta-lactamase
penam
antibiotic inactivation
106 3002681
chloramphenicol acetyltransferase (CAT)
phenicol antibiotic
antibiotic inactivation
110 3002565
AAC(6’)
aminoglycoside antibiotic
antibiotic inactivation
114 3003012
trimethoprim resistant dihydrofolate reductase dfr
diaminopyrimidine antibiotic
antibiotic target replacement
116 3002733
quinolone resistance protein (qnr)
fluoroquinolone antibiotic
antibiotic target protection
120 3002604
ANT(3”)
aminoglycoside antibiotic
antibiotic inactivation
124 3002263
IND beta-lactamase
carbapenem
antibiotic inactivation
128 3002828
Miscellaneous ABC-F subfamily ATP-binding cassette ribosomal protection proteins
macrolide antibiotic
antibiotic target protection
129 3000198
fosfomycin thiol transferase
phosphonic acid antibiotic
antibiotic inactivation
130 3001972
CTX-M beta-lactamase
cephalosporin
antibiotic inactivation
@@
Appendix D Baselines
@@
Appendix E Evaluation Metrics.
To fully evaluate the model performance, we used several evaluation metrics including Accuracy, Macro Pre, Macro Recall, Macro F1 and Weighted F1.
Precision: indicates the precision rate, which is the ratio of the number of true positive samples to the number of samples classified as positive. precision = TP / (TP + FP)
Recall: the ratio of the number of true positive samples to the number of all positive samples in the sample set, Recall = TP / (TP + FN)
F1score: used to combine Precision and Recall, F1-Score = 2 * Precision * Recall / (Precision + Recall)
Appendix F Overview of Downstream Tasks
Operon Prediction Task is to identify the transcription factor binding sites that have the strongest correlation with operon regulation in the gene regulatory network [9, 13], which helps us to understand the mechanism and network of gene regulation and reveals the key interactions between transcription factors and operons [46]. In general, operons are composed of a set of multiple genes close to each other.
@@ Appendix F Overview of Downstream Tasks
Overview of Pathogens. Pathogenic bacteria are a group of bacteria that can cause disease and can pose a significant threat to human health compared to commensal non-pathogenic bacteria in the human body with their ability to invade the host and cause disease [27]. The existing methods are limited to long genome sequences and are dominated by short sequence (¡3 kb) inputs, which are unable to directly process long genome sequences at the million bp level. Hence, the development of accurate deep learning algorithms for the precise identification of pathogens (causative agents) can help in the development of new therapies and vaccines, and improve the ability to respond to the threat of drug resistance.
@@
Overview. The Nitrogen cycle is a collection of important biogeochemical pathways in the Earth’s ecosystems, and quantitatively studying the functional genes related to the N cycle and linking them to environmental and ecological processes is one of the important focuses in environmental genomics research [16]. Currently, macrogenome sequencing has been widely used to characterize gene families involved in N cycling processes. However, existing methods usually face the problems of inefficient database searches, unclear direct lineage classification, and low coverage of N cycle genes and/or gene (sub)families when analyzing N cycle gene families in macrogenomes. Consequently, accurate prediction of the functional categories of genes or proteins related to nitrogen metabolism in microorganisms is important for analyzing important resources in the process of nitrogen cycling and understanding the nitrogen cycle in ecosystems.
Appendix G Detailed Classification Results on CARD Dataset

### v2

@@ 4.2 Ablation Study
Table 6: Ablation Study of M.F1 on CARD-D.
@@
Figure 3: Ablation studies of Our Proposed Modules on Four Downstream Tasks.
Ablation Study on the Performance of MGM and TMC. We perform the ablation study to investigate the effectiveness of our proposed components. Tab. 6 compares the performance of FGBERT without MGM and TMC module across four datasets. The decrease in M.F1 score after removing the MGM and TMC, respectively, highlights their roles in enhancing model performance across different tasks. While TMC contributes to performance, it is evident that MGM has a more substantial impact. Fig. 3 illustrates four more metrics on the CARD dataset.
Ablation Study on Visualization Results of MGM. We conduct a visualization experiment to validate the effectiveness of MGM in the One-to-Many scenario. ATP synthases can exhibit different functions in different organisms to adapt to their respective environmental conditions, even though the basic functions are the same [22].
@@
Table 7: T-SNE Visualization of Different Embeddings for ATP Synthases. Each dot denotes a sequence and is colored according to different functions.
@@
We collect 1177 ATP synthase sequences from UniProt [1] and color them according to six taxonomies, i.e. Hydrolases (35.6%, dark green), Isomerases (3.6%, orange), Ligases (35.7%, blue), Lyases (10.7%, pink), Oxidoreduct (8.9%, light green) and Transferases (5.3%, yellow) (numbers in parentheses indicate category ratios). Fig.7 (a) shows the clustering results of ATPase protein embeddings without genome contextual analysis, indicating that more dispersed clustering results of genes in different genome contexts. Fig.7 (b) presents the clustering results of ATPase embeddings after our designed protein-based gene embedding, which shows genes belonging to the same category (the same color) apparently cluster together. Concretely, Isomerases (3.5%, orange) and Transferases (5.3%, yellow), which account for the smallest percentage, cluster together, whereas, in the left plot, these two are scattered. This demonstrates that our proposed MGM can resolve the One-to-Many problem between sequences and functions.
Table 8: Clustering Results of TMC for Many-to-One problem.
Method NMI ARI
Silhouette
Coefficient
+Tokenizer 0.44 0.34 0.7
+MGM 0.66 0.51 0.72
+TMC 0.72 0.59 0.75
Ablation Study on Clustering results of TMC. We further explore the impact of integrating TMC into our model on gene operon prediction. Initially, we evaluate FGBERT without TMC module. Afterwards, we add TMC module, and K-means are performed using the embeddings from the network’s intermediate layers. We compute the Normalized Mutual Information (NMI), Adjusted Rand Index (ARI), and Silhouette Coefficient Index to measure the clustering quality. The results in Tab. 8 show significant improvements in ARI and Silhouette Coefficient with TMC, highlighting its effectiveness in enhancing clustering performance across various metrics.
Ablation Study on Context-Aware Tokenizer. We replace the FGBERT’s tokenizer with ESM2 and BPE representations and assessed performance across downstream tasks. Tab. 9 shows that FGBERT outperforms both methods in capturing complex gene sequence-function relationships. Although DNABERT2 compresses sequences effectively with BPE, our context-aware tokenizer demonstrates superior accuracy in metagenomic data analysis. For complete results, see Appendix Tab. 17.
Table 9: The ablation study using protein-based gene representation as a context-aware tokenizer.
Dataset Method Acc M.Pre M.Recall M.F1 W.Pre W.Recall W.F1
E-K12 FGBERT 0.68 0.68 0.61 0.61 0.77 0.67 0.65
ESM2 0.54 0.51 0.48 0.49 0.61 0.51 0.52
BPE 0.59 0.58 0.55 0.54 0.67 0.54 0.55
CARD-A FGBERT 0.91 0.77 0.8 0.78 0.9 0.91 0.9
ESM2 0.82 0.74 0.76 0.73 0.87 0.82 0.81
BPE 0.84 0.75 0.77 0.75 0.88 0.86 0.86
4.3 Model Efficiency Analysis
Table 10: Comparative Analysis on Tokenization Efficiency: Time(s) vs. Memory (MB). Each point denotes a specific dataset, with the size indicating its scale.
@@ 4.3 Model Efficiency Analysis
We analyze the time complexity and memory efficiency of our tokenizer compared to four genomic pre-trained methods on six datasets in Fig. 10. Our tokenizer demonstrates superior efficiency, achieving a significant reduction in both time and memory usage. Notably, on NCyc dataset (brown) with 213,501 sequences, ours reduces processing time by 31.05% and memory usage by 94.33% compared to DNABERT2. For the CARD dataset (orange) with 1,966 sequences, time and memory usage decreased by 61.70% and 58.53%. Although HyenaDNA uses less memory on Operons, CARD, VFDB, and ENZYME datasets, it underperforms ours in time cost and overall performance.
Table 11: The Scalability of FGBERT.
Model Pre-training Data # Layers Hidden Dim # Heads Operons (500/1000/1500 ep) CARD-A (500/1000/1500 ep) CARD-D (500/1000/1500 ep) CARD-R (500/1000/1500 ep) FGBERT-T 50M 10 640 5 50.9 / 61.5 / 63.5 74.6 / 88.6 / 90.7 72.4 / 83.7 / 86.4 77.3 / 90.1 / 91.1 FGBERT-S 100M 19 1280 10 55.1 / 65.4 / 65.9 80.4 / 90.1 / 91.2 76.9 / 85.2 / 87.5 81.4 / 91.4 / 93.0 FGBERT-B 150M 25 2560 25 57.1 / 66.7 / 67.6 82.7 / 91.1 / 93.2 81 / 87.6 / 90.0 83.7 / 93.4 / 94.8
To further explore model scalability, we train three FGBERT variants—FGBERT-T, FGBERT-S, and FGBERT-B—differing in layers, hidden dimensions, and attention heads. Each variant is trained on 50 million, 100 million, and 150 million sequences to assess the impact on time and memory during pre-training, as shown in Tab. 11.
@@ 4.4 Sensitivity Analysis
Table 12: Sensitivity w.r.t Hyper-parameters $\alpha,b$ of CARD dataset on AMR Gene Family.
Our sensitivity analysis indicates that FGBERT can be optimized effectively using small batch size $b$ without larger performance degradation shown in Fig. 12. This is important for resource-constrained scenarios, highlighting our model maintains good performance even with limited data. We choose a batch size of 1000. Next, we analyze another hyper-parameter of balance ratio $\alpha$ on CARD dataset. FGBERT obtains good performance for different values of $\alpha$, demonstrating its robustness and insensitivity to this hyper-parameter. We set $\alpha$ to be 0.4.
4.5 Model Configurations.
All our experiments are performed on 4 NVIDIA V100 GPUs and the PyTorch framework. The encoder of FGBERT is initialized with Roberta [11]. For pre-training, the parameters are set as follows: batch size of 1000, 19 encoder layers, 10 attention heads, an embedding dimension of 1280, and relative position encoding. During the pre-training stage, the model is trained for 500 epochs using the AdamW optimizer [39] with a weight decay of 0.02. The learning rate starts at 1e-5, warms up to 1e-4 over the first 5000 steps, and then decreases back to 1e-5 following a cosine decay schedule. The complete model consists of 954.73 million parameters and requires 2.55 billion FLOPs, as detailed in Appendix Tab. 15.
@@ 5 Conclusion
In this paper, we propose a new idea of protein-based gene representation, preserving essential biological characteristics within each gene sequence. With the new context-aware tokenizer, we propose MGM, a gene group-level pre-training task, designed to learn the interactions between genes. Additionally, we develop TMC, a contrastive learning module to generate multiple positive and negative samples to distinguish the gene sequences. MGM and TMC constitute a joint pre-training model, FGBERT for metagenomic data. Our experiments and visualizations demonstrate the superior performance of our model. For the future, it remains to be explored how to incorporate multi-omics data, such as metabolomics, into our metagenomic pre-trained model.
Appendix A Appendix / supplemental material
Optionally include supplemental material (complete proofs, additional experiments and plots) in appendix. All such materials SHOULD be included in the main submission.
Appendix B Methods.
B.1 Negative Sample Sampling Strategy.
The selection of negative samples is mapped based on the Euclidean distance between EC numbers. Each EC number is considered a unique functional category or cluster, with seven clusters corresponding to seven distinct EC numbers. For each anchor sample, ten negative samples are chosen based on their proximity within the distance mapping, selecting the nearest in terms of Euclidean distance. To provide a concise overview, the clusters and their corresponding entries are summarized in Tab. 13.
Table 13: The specific details of the negative sample sampling strategy.
EC Entry EC number
1 F2JXJ3 1.4.3.20
2 Q3TTA7 2.3.2.27
3 Q8K337 3.1.3.36
4 Q9LZI2 4.1.1.35
5 Q79FW5 5.4.2.12
6 Q9HN83 6.1.1.2
7 Q33821 7.1.1.2
Appendix C Implementation Details.
C.1 Dataset.
MGnify dataset includes genome sequence information from different microbial communities and their Biome Names (e.g., "Engineered", "Host-associated", and "Control"), as well as their corresponding Sample Classes, Sample Numbers, and Spectrum Categories. Lineage Category refers to the taxonomic information of microorganisms, which describes the taxonomic hierarchy of microorganisms from Domain (e.g., ”root” or ”Engineered”) to Species (e.g., ”Continuous culture” or ”Saliva”), which can be observed in Tab. 14.
Table 14: Microbial Species and Genomic Data Across Various Environments in the MGnify Dataset. Each column is catalogues, species, and genomes, respectively.
@@ Appendix A Appendix / supplemental material
Table 15: Examples of Operons Dataset.
GeneID Operon name Description of Operon name
@@
2973 qseB DNA-binding transcriptional activator QseB qseBC None
2980 cpdA cAMP phosphodiesterase nudF-yqiB-cpdA-yqiA-parE None
2985 ygiC putative acid–amine ligase YgiC tolC-ygiBC tolC-ygiBC
@@
Model ID ARO Accession AMR Gene Family Drug Class Resistance Mechanism
1 3002999 CblA beta-lactamase cephalosporin antibiotic inactivation
5 3002867 trimethoprim resistant dihydrofolate reductase dfr diaminopyrimidine antibiotic antibiotic target replacement
7 3001989 CTX-M beta-lactamase cephalosporin antibiotic inactivation
10 3002244 CARB beta-lactamase penam antibiotic inactivation
20 3002012 CMY beta-lactamase cephamycin antibiotic inactivation
@@
E.coil K12 RegulonDB Tab. 15 lists information such as operon names, descriptions, and names of genes included for some of the EK12 datasets.
CARD Tab. 15 shows a partial list of instances from the CARD dataset, including data IDs, ARO (Antibiotic Resistance Ontology) Accession IDs, AMR (Antimicrobial Resistance) category names, drug category names, and Resistance Mechanism names.
C.2 Baselines
@@
C.3 Model Configurations.
All our experiments are performed on 4 NVIDIA V100 GPUs and the PyTorch framework. The encoder of FGBERT is initialized with Roberta [11]. The parameters for pre-training: batch size = 1000, 19 encoder layers, 10 attention heads, embedding dimension = 1280, and relative position encoding. During the pre-training stage, the model is trained for 500 epochs. We use the AdamW optimizer [39] with a weight decay of 0.02. The learning rate is initialized to 1e-5 and warmed up to 1e-4 after 5000 epochs. Then, it is reduced to 1e-5 by a cosine decay strategy. The overall model comprises 954.73M parameters with a computational load of 2.55B FLOPs, detailed in Tab. 15.
Table 15: The parameters and network layers learned by pre-training.
Batch Size 1000
Encoder Layer 19
Attention Heads 10
Embedding Dimension 1280
Positive Encoding Relative Positive Encoding
Epochs 1000
Optimizer AdamW
Weight Decay 0.02
Learning Rate 1e-5→1e-4
Learning Rate Schedule Cosine
Warm-up Steps 5000
C.4 Overview of Downstream Tasks
Operon Prediction Task is to identify the transcription factor binding sites that have the strongest correlation with operon regulation in the gene regulatory network [9, 13], which helps us to understand the mechanism and network of gene regulation and reveals the key interactions between transcription factors and operons [51]. Operons are generally composed of a set of multiple genes close to each other.
@@ C.4 Overview of Downstream Tasks
Overview of Pathogens. Pathogenic bacteria are a group of bacteria that can cause disease and can pose a significant threat to human health compared to commensal non-pathogenic bacteria in the human body with their ability to invade the host and cause disease [28]. The existing methods are limited to long genome sequences and are dominated by short sequence (<3 kb) inputs, which cannot process long genome sequences at the million bp level directly. Hence, the development of accurate deep learning algorithms for the precise identification of pathogens (causative agents) can help in the development of new therapies and vaccines and improve the ability to respond to the threat of drug resistance.
@@
Overview. The Nitrogen cycle is a collection of important biogeochemical pathways in the Earth’s ecosystems, and quantitatively studying the functional genes related to the N cycle and linking them to environmental and ecological processes is one of the important focuses in environmental genomics research [16]. Currently, metagenome sequencing has been widely used to characterize gene families involved in N cycling processes. However, existing methods usually face the problems of inefficient database searches, unclear direct lineage classification, and low coverage of N cycle genes and/or gene (sub)families when analyzing N cycle gene families in metagenomes. Consequently, accurate prediction of the functional categories of genes or proteins related to nitrogen metabolism in microorganisms is important for analyzing important resources in the process of nitrogen cycling and understanding the nitrogen cycle in ecosystems.
Appendix D Experimental Results.
D.1 More Detailed Ablation Experiments
The results of Tab. 16 demonstrate each module’s significant role in the overall performance.
Table 16: The result of FGBERT across eight downstream tasks involving seven evaluation metrics.
Dataset Method Acc M.Pre M.Recall M.F1 W.Pre W.Recall W.F1 E-K12 FGBERT 0.68 0.68 0.61 0.61 0.77 0.67 0.65 w/o. MGM 0.58(↓14.71%) 0.56(↓17.65%) 0.53(↓13.11%) 0.50(↓18.03%) 0.58(↓24.68%) 0.57(↓14.93%) 0.56(↓13.85%) w/o. Triplet 0.61(↓10.29%) 0.60(↓11.76%) 0.56(↓8.20%) 0.54(↓11.48%) 0.61(↓20.78%) 0.63(↓5.97%) 0.59(↓9.23%) CARD-A FGBERT 0.91 0.77 0.8 0.78 0.9 0.91 0.9 w/o. MGM 0.80(↓12.09%) 0.73(↓5.19%) 0.76(↓5.00%) 0.72(↓7.69%) 0.87(↓3.33%) 0.80(↓12.09%) 0.79(↓12.22%) w/o. Triplet 0.87(↓4.40%) 0.72(↓6.49%) 0.75(↓6.25%) 0.73(↓6.41%) 0.86(↓4.44%) 0.87(↓4.40%) 0.88(↓2.22%) CARD-D FGBERT 0.85 0.56 0.6 0.57 0.85 0.85 0.85 w/o. MGM 0.80(↓5.88%) 0.48(↓14.29%) 0.50(↓16.67%) 0.47(↓17.54%) 0.81(↓4.71%) 0.80(↓5.88%) 0.81(↓4.71%) w/o. Triplet 0.82(↓3.53%) 0.49(↓12.50%) 0.58(↓3.33%) 0.52(↓8.77%) 0.82(↓3.53%) 0.82(↓3.53%) 0.82(↓3.53%) CARD-R FGBERT 0.92 0.75 0.67 0.69 0.91 0.92 0.91 w/o. MGM 0.88(↓4.35%) 0.65(↓13.33%) 0.64(↓4.48%) 0.63(↓8.70%) 0.89(↓2.20%) 0.90(↓2.17%) 0.89(↓2.20%) w/o. Triplet 0.90(↓2.17%) 0.67(↓10.67%) 0.65(↓2.99%) 0.66(↓4.35%) 0.89(↓2.20%) 0.9(↓2.17%) 0.89(↓2.20%) VFDB FGBERT 0.91 0.83 0.71 0.75 0.91 0.91 0.91 w/o. MGM 0.85(↓6.59%) 0.7(↓15.66%) 0.58(↓18.31%) 0.60(↓20.00%) 0.85(↓6.59%) 0.84(↓7.69%) 0.85(↓6.59%) w/o. Triplet 0.87(↓4.40%) 0.81(↓2.41%) 0.66(↓7.04%) 0.70(↓6.67%) 0.88(↓3.30%) 0.89(↓2.20%) 0.88(↓3.30%) ENZYME FGBERT 0.99 0.99 0.98 0.99 0.99 0.98 0.99 w/o. MGM 0.95(↓4.04%) 0.96(↓3.03%) 0.95(↓3.06%) 0.96(↓3.03%) 0.95(↓4.04%) 0.95(↓3.06%) 0.95(↓4.04%) w/o. Triplet 0.95(↓4.04%) 0.97(↓2.02%) 0.96(↓2.04%) 0.96(↓3.03%) 0.97(↓%2.02) 0.96(↓2.04%) 0.96(↓3.03%) PATRIC FGBERT 0.95 0.63 0.68 0.64 0.92 0.94 0.93 w/o. MGM 0.63(↓33.68%) 0.34(↓46.03%) 0.28(↓58.82%) 0.30(↓53.13%) 0.67(↓27.17%) 0.63(↓32.98%) 0.64(↓31.18%) w/o. Triplet 0.83(↓12.63%) 0.58(↓7.94%) 0.57(↓16.18%) 0.56(↓12.50%) 0.81(↓11.96%) 0.83(↓11.70%) 0.81(↓12.90%) NCycDB FGBERT 0.99 0.99 0.99 0.99 0.99 0.99 0.99 w/o. MGM 0.92(↓7.07%) 0.94(↓5.05%) 0.95(↓4.04%) 0.91(↓8.08%) 0.93(↓6.06%) 0.95(↓4.04%) 0.97(↓2.02%) w/o. Triplet 0.95(↓4.04%) 0.96(↓3.03%) 0.96(↓3.03%) 0.95(↓4.04%) 0.97(↓2.02%) 0.96(↓3.03%) 0.97(↓2.02%)
Table 17: The ablation study using protein-based gene representation as a context-aware tokenizer.
Dataset Method Acc M.Pre M.Recall M.F1 W.Pre W.Recall W.F1
E-K12 FGBERT 0.68 0.68 0.61 0.61 0.77 0.67 0.65
ESM2 0.54 0.51 0.48 0.49 0.61 0.51 0.52
BPE 0.59 0.58 0.55 0.54 0.67 0.54 0.55
CARD-A FGBERT 0.91 0.77 0.8 0.78 0.9 0.91 0.9
ESM2 0.82 0.74 0.76 0.73 0.87 0.82 0.81
BPE 0.84 0.75 0.77 0.75 0.88 0.86 0.86
CARD-D FGBERT 0.85 0.56 0.6 0.57 0.85 0.85 0.85
ESM2 0.82 0.49 0.58 0.52 0.83 0.82 0.82
BPE 0.82 0.49 0.58 0.52 0.83 0.82 0.82
CARD-R FGBERT 0.92 0.75 0.67 0.69 0.91 0.92 0.91
ESM2 0.88 0.67 0.66 0.67 0.87 0.88 0.89
BPE 0.9 0.69 0.67 0.68 0.9 0.9 0.9
VFDB FGBERT 0.91 0.83 0.71 0.75 0.91 0.91 0.9
ESM2 0.85 0.69 0.57 0.6 0.85 0.85 0.84
BPE 0.86 0.7 0.56 0.59 0.86 0.86 0.85
ENZYME FGBERT 0.99 0.99 0.98 0.99 0.99 0.98 0.98
ESM2 0.95 0.96 0.95 0.96 0.95 0.95 0.95
BPE 0.96 0.96 0.95 0.96 0.96 0.96 0.96
PATRIC FGBERT 0.95 0.63 0.68 0.99 0.92 0.94 0.99
ESM2 0.63 0.34 0.27 0.3 0.63 0.61 0.62
BPE 0.56 0.28 0.28 0.27 0.6 0.56 0.57
NCycDB FGBERT 0.99 0.99 0.99 0.99 0.99 0.99 0.99
ESM2 0.56 0.28 0.28 0.27 0.6 0.56 0.57
BPE 0.63 0.33 0.27 0.29 0.66 0.62 0.63
We have replaced the original tokenizer with ESM2 representations and BPE and assessed its impact on various downstream tasks. Tab. 17 demonstrates that the FGBERT tokenizer outperforms ESM2 and BPE representations across various metrics, indicating its efficacy in capturing intricate relationships between gene sequences and their functions. While DNABERT-2’s use of BPE efficiently compresses sequences, our context-aware tokenizer excels in precision with metagenomic data.
D.2 Detailed Classification Results on CARD Dataset
This section provides more detailed ablation and analytical experiments on the individual modules.